OpenStudy (anonymous):

water is running out a conical funnel at the rate of 1cu in per sec. if the radius of the base of the funnel is 4an and the altitude is 8in. find the rate at which the water level is dropping when it is 2 in from the top.

OpenStudy (amistre64):

change of rates eh....

OpenStudy (amistre64):

v' = 1 ft^3 per sec; or simply 1 the rate of h is what we wanna determine

OpenStudy (amistre64):

volume of a cone = (1/3)(base area)(height)

OpenStudy (amistre64):

the relation of height to raduis is given by the cross section of the cone

OpenStudy (amistre64):

h = -2r+8 right?

OpenStudy (anonymous):

can you show the solution?

OpenStudy (amistre64):

v = (1/3)(pi 2^2)(-2r +8)

OpenStudy (amistre64):

im working on it; havent got to the end yet lol

OpenStudy (amistre64):

pi 2^2 was meant to be pi r^2

OpenStudy (amistre64):

lets re work that with h as the variabe so we can see the rate of change with respect to h and not r

OpenStudy (amistre64):

r = (h-8)/-2 or (8-h)/2

OpenStudy (amistre64):

v = (1/3)(pi ((8-h)/2)^2)(h) perhaps?

OpenStudy (amistre64):

\[v = \frac{h}{3}*\frac{\pi.(8-h)^2}{4}\]

OpenStudy (amistre64):

64hpi +h^3pi -16h^2pi --------------------- = v derive now 12

OpenStudy (amistre64):

64pi +3h^2pi -32hpi --------------------- = v' derive now 12

OpenStudy (amistre64):

pi(64 +3h^2 -32h) = 12 64 +3h^2 -32h = 12/pi right?

OpenStudy (anonymous):

ang gulo wah

OpenStudy (amistre64):

maybe i mess that up alittle dv/dh = that up there dv/dt = 1 and dv/dh dh/dt = dv/dt

OpenStudy (amistre64):

we wanna find dh/dt sooo dh/dt = dh/dv dv/dt ; gotta take the inverse if thats gonna work

OpenStudy (amistre64):

dh/dt is the inverse of dh/dv right :)

OpenStudy (amistre64):

on paper i get dh/dt = -1/(9pi)